OpenStudy (anonymous):

Log4 (x+1) + log4 (x-3) = log4 12 solve for x

OpenStudy (sirm3d):

use the property of Log \[\Large \log_b M+\log_b N = \log_b {MN}\]

OpenStudy (anonymous):

I don't know what to do with that

OpenStudy (sirm3d):

your b =4 , M = (x+1), N = (x-3)

OpenStudy (sirm3d):

\[ \Large{\log_b(M) \quad\;\;\;\; + \log_b(N)\quad \quad\quad=\log_b(MN)\\\log_4(x+1) + \log_4(x-3)}\;\;\;\;= \]

OpenStudy (anonymous):

Then what?

OpenStudy (sirm3d):

what do you have so far?

OpenStudy (anonymous):

The same equation

OpenStudy (sirm3d):

log4[(x+1)(x-3) = log4 12

OpenStudy (anonymous):

Do you distribute?

OpenStudy (sirm3d):

FOIL (x+1)(x-3)

OpenStudy (anonymous):

What happens next?